OpenStudy (anonymous):

Solve for t e^(2t+1)=9e^(1-t)

OpenStudy (solomonzelman):

\(\large\color{slate}{ e^{2t+1}=9e^{1-t} }\)

OpenStudy (anonymous):

i took the natural log of both sides and sort of solved from there but my answers were .2 off so im not sure i did it right

OpenStudy (solomonzelman):

yes, you can do something very similar to what you have done as well.

OpenStudy (solomonzelman):

For some reason I thought of, \(\large\color{slate}{ e^{2t+1}=\color{blue}{9}e^{1-t} }\) \(\large\color{slate}{ e^{2t+1}=\color{blue}{e^{\ln 9}}e^{1-t} }\) \(\large\color{slate}{ e^{2t+1}=e^{\ln 9+1-t} }\)

OpenStudy (anonymous):

hmmm ok. I wouldnt have immediately thought of that but I understand it

OpenStudy (anonymous):

Solomon, your final answer for t was around .732, correct?

OpenStudy (amoodarya):

sorry , i write another way to solve \[e^{2t+1}=9e^{1-t}\\e^{2t}*e=9*e*e^{-t}\\e^{2t}=9*e^{-t}\\a=e^t\\so\\a^2=9*\frac{1}{a}\\a^3=9\]

OpenStudy (solomonzelman):

idk, about the approximation, but: \(\large\color{slate}{ 2t+1=\ln(9)+1-t }\) \(\large\color{slate}{ 3t+1=\ln(9)+1 }\) \(\large\color{slate}{ 3t=\ln(9) }\) \(\large\color{slate}{ t=\frac{1}{3}\ln(9) }\)

OpenStudy (anonymous):

Yes thats what I got, Solomon. Thank you.

OpenStudy (solomonzelman):

yw